OpenStudy (babybritt002):

Iron is sold in sheets that are 2.0 cm thick and 120.0 cm wide. You want to buy a sheet of iron with a mass of 50,000.0 g. What must be the length of the sheet of iron?

OpenStudy (sleepyhead314):

hmm are you told any more information? like the weight of the 2.0cm by 120.0 cm sheet?

OpenStudy (sleepyhead314):

http://www.answers.com/Q/What_is_the_weight_of_1_cubic_meter_of_iron 1 m^3 => 7,874 kg we should use this ratio to find your answer

OpenStudy (sleepyhead314):

1 m^3 = 1000000 cm^3

OpenStudy (sleepyhead314):

7874 kg = 7874000 g

OpenStudy (sleepyhead314):

this ratio simplifies to 500 cm^3 = 3937 g of iron

OpenStudy (sleepyhead314):

volume of the sheet would be 2 * 120 * x = 240*x cm^3

OpenStudy (sleepyhead314):

we can match ratio to ratio to find "x" 240*x 50000 --------- = ---------- 500 3937

OpenStudy (sleepyhead314):

http://www.wolframalpha.com/input/?i=240x%2F500+%3D+50000%2F3937 scroll down to "solution" and click "approximate form"